# A relationship between the Kauffman bracket skein algebras and Roger-Yang skein algebras of some small surfaces

### v1

@@ 6 Classification of Representations of $\mathcal{S}^{RY}(\Sigma_{0,2,2})$
For the remainder of this section we explore to what extent we can recover $\rho$ from its classical shadow data. We adapt the method of [36] to show that when the classical shadow data satisfy some polynomial conditions, there is a unique finite-dimensional irreducible representation of $\mathcal{S}^{RY}(\Sigma_{0,2,2})$ with that classical shadow data.
@@
Let $A$ be a primitive $N$th root of 1-1 (for $N$ odd). Let $(t_{1},t_{2},t_{3},d_{0},d_{1})\in\mathbb{C}^{5}$ such that
@@
By assumption, $\rho(\gamma)\neq 0$, so one of its eigenspaces, say $V_{k_{0}}$ is nonzero. The map $U_{k_{0}+N-1}\cdots U_{k_{0}}:V_{k_{0}}\to V_{k_{0}}$ is nonzero by Lemma 6.5. Therefore, it has an eigenvalue $u\neq 0$ with associated eigenvector $\mathbf{v}_{k_{0}}$. Set $\mathbf{v}_{1}=U_{N}U_{N-1}\cdots U_{k_{0}}\mathbf{v}_{k_{0}}$. By construction, $U_{k_{0}+N-1}\cdots U_{k_{0}}\mathbf{v}_{k_{0}}=u\mathbf{v}_{k_{0}}$, so multiplying by $U_{N}U_{N-1}\cdots U_{k_{0}}$ we obtain
@@
The left hand side is a polynomial in $\sqrt{v_{1}v_{2}}U$ and $\sqrt{v_{1}v_{2}}D$ over $\mathbb{C}[v_{1},v_{2}]$. Consider a monomial of length $m$, where $n$ of the terms are $\sqrt{v_{1}v_{2}}U$ and mnm-n are $\sqrt{v_{1}v_{2}}D$. It is a property of $T_{N}$ that all the monomials will be of odd degree, so $m$ is odd. The monomial sends $V_{k}$ to $V_{k+n-(m-n)}=V_{k+2n-m}$. As $T_{N}(\sqrt{v_{1}v_{2}}(U+D))=t_{1}\text{id}_{V}$ is homothety, it must fix all the $V_{k}$. Hence, the only monomials with nonzero coefficient must satisfy $2n-m\equiv 0\text{ mod }N$. As $n,m\in\{0,\ldots,N\}$, we have $(m,n)$ is $(N,0)$ or $(N,N)$. As $T_{N}(\sqrt{v_{1}v_{2}}(U+D))$ has degree $N$, $T_{N}(\sqrt{v_{1}v_{2}}(U+D))=\sqrt{v_{1}v_{2}}^{N}U^{N}+\sqrt{v_{1}v_{2}}^{N}D^{N}$.
@@
Equations 17 and 18 have no common roots $x$, so they cannot hold simultaneously. Hence, we must have $d_{0}+d_{1}=0$. ∎
@@
We may think of $d_{0}+d_{1}=0$ and $T_{N}(2-d_{0}^{2})-2=-t_{1}^{2}-t_{2}^{2}-t_{3}^{2}-t_{1}t_{2}t_{3}$ as analogues of the puncture invariant conditions for the skein algebra of the torus by [36, Lem. 7].
@@
Note that if we substitute $d_{0}=A+A^{-1}$ and $d_{1}=-A-A^{-1}$, $P_{k}$ becomes $A^{2}-A^{-2}$, meaning that it corresponds under the homomorphism $\varphi:\mathcal{S}^{RY}(\Sigma_{0,2,2})\to\mathcal{S}(\Sigma_{1,0,0})$ to a nullhomotopic loop in the closed torus. Similarly, if we substitute $d_{0}=A+A^{-1}$ into (19), we recover the condition on the puncture invariant for the once punctured torus at $p=-A^{2}-A^{-2}$, namely, $T_{N}(-A^{2}-A^{-2})=-t_{1}t_{2}t_{3}-t_{1}^{2}-t_{2}^{2}-t_{3}^{3}+2$, given in [36, Lem. 7]. Our investigation was motivated by whether all of the representations of $\mathcal{S}^{RY}(\Sigma_{0,2,2})$ were constructed this way, which we found to the negative.
Problem 6.12.
@@ 7.1 A geometric basis for $S^{RY}(\Sigma_{0,2,2})$
First consider the case where $\left[\begin{smallmatrix}n_{1}\\ k_{1}\end{smallmatrix}\right]_{g}$ and $\left[\begin{smallmatrix}n_{2}\\ k_{2}\end{smallmatrix}\right]_{g}$ are both loops. We represent them as loops in the four-punctured sphere, respectively $(\frac{n_{1}+k_{1}}{2},n_{1})$ and $(\frac{n_{2}+k_{2}}{2},n_{2})$. Observe we must have $\text{gcd}(\frac{n_{i}+k_{i}}{2},n_{i})=1$, and recall by [2, Lemma 4.1] (see also Section 2.2.5 of [18]) that the geometric intersection number of loops in the pillowcase $(N_{1},K_{1})$ and $(N_{2},K_{2})$ is $2|N_{1}K_{2}-N_{2}K_{1}|$ when $\text{gcd}(N_{i},K_{i})=1$.
Now suppose one of the curves, say $\left[\begin{smallmatrix}n_{1}\\ k_{1}\end{smallmatrix}\right]_{g}:=\theta$ is an arc, and the other $\left[\begin{smallmatrix}n_{2}\\ k_{2}\end{smallmatrix}\right]_{g}:=\omega$ is a loop, then let $\theta^{\prime}=v_{1}v_{2}\left[\begin{smallmatrix}n_{1}\\ k_{1}\end{smallmatrix}\right]_{g}^{2}-2=\left[\begin{smallmatrix}2n_{1}\\ 2k_{1}\end{smallmatrix}\right]_{g}$, which is the boundary of a regular neighborhood of $\theta$, as in Example 3.2. Then, from Figure 4 (Left), we see the intersection number of $\omega$ and $\theta^{\prime}$ is twice that of $\theta$ and $\omega$, and, as $\omega$ and $\theta^{\prime}$ are loops. One can show that if $\left[\begin{smallmatrix}n_{1}\\ k_{1}\end{smallmatrix}\right]_{g}$ is an arc then $\text{gcd}(n_{1}+k_{1},2n_{1})=1$, so we can apply [2, Lemma 4.1] as before.
@@ 7.2 Description of pre-image of torus knots
We consider only the case where $0<K<N$, as the other case is similar. By [19, Lemma 1], for any $(N,K)$ such that $N\geq 3$, $0<K<N$ and $\text{gcd}(N,K)=1$, there exist $u,v,w,z$ such that $uz-vw=\pm 1$, $u+w=N$, $v+z=K$, $w\in\{1,\ldots,N-1\}$, $u\in\{1,\ldots,N-2\}$, and $v,z\in\mathbb{Z}_{>0}$. It follows that the pairs $(u,v),(w,z),$ and (uw,zv)(u-w,z-v) are each relatively prime.
@@ 7.3 Structural constants for infinite families of products
To simply the formulas, we use normalized Chebyshev polynomials $\overline{T}_{n}(x)$, defined by $\overline{T}_{0}(x)=1$ and for all $n\geq 1$, $\overline{T}_{n}(x)=T_{n}(x)$.
@@
$b(p,k)$ $=\begin{cases}A^{-k}\left(\delta_{[p-k]}^{-1}\frac{p-k+1}{4}+\displaystyle{\sum_{h=1}^{\frac{p-k+[p-k]-2}{4}}}h([p-k-4h+2]_{A}+[p-k-4h]_{A})\right)&k\leq p-3\quad\text{and}\quad k\equiv p-3\mod 2\\ 0&\text{otherwise}\end{cases}$

### v2

@@ 6.2 Explicit representations of $\mathcal{S}^{RY}(\Sigma_{0,2,2})$
For the remainder of this section we explore to what extent we can recover $\rho$ from its classical shadow data. We adapt the method of [41] for the closed torus and [26] for $U_{q}(so_{3})$ to show that when the classical shadow data satisfy some polynomial conditions, there is a unique finite-dimensional irreducible representation of $\mathcal{S}^{RY}(\Sigma_{0,2,2})$ with that classical shadow data.
@@
Let $A$ be a primitive $2N$th root of unity for $N$ odd. Let $(t_{1},t_{2},t_{3},d_{0},d_{1})\in\mathbb{C}^{5}$ such that
@@
By assumption, $\rho(\gamma)\neq 0$, so one of its eigenspaces, say $V_{k_{0}}$, is nonzero. The map $U_{k_{0}+N-1}\cdots U_{k_{0}}:V_{k_{0}}\to V_{k_{0}}$ is nonzero by Lemma 6.7. Therefore, it has an eigenvalue $u\neq 0$ with associated eigenvector $\mathbf{v}_{k_{0}}$. Set $\mathbf{v}_{1}=U_{N}U_{N-1}\cdots U_{k_{0}}\mathbf{v}_{k_{0}}$. By construction, $U_{k_{0}+N-1}\cdots U_{k_{0}}\mathbf{v}_{k_{0}}=u\mathbf{v}_{k_{0}}$, so multiplying by $U_{N}U_{N-1}\cdots U_{k_{0}}$ we obtain
@@
The left hand side is a polynomial in $\sqrt{v_{1}v_{2}}U$ and $\sqrt{v_{1}v_{2}}D$ over $\mathbb{C}[v_{1},v_{2}]$. Consider a monomial of length $m$, where $n$ of the terms are $\sqrt{v_{1}v_{2}}U$ and mnm-n are $\sqrt{v_{1}v_{2}}D$. It is a property of $T_{N}$ that all the monomials are of odd degree, so $m$ is odd. The monomial sends $V_{k}$ to $V_{k+n-(m-n)}=V_{k+2n-m}$. As $T_{N}(\sqrt{v_{1}v_{2}}(U+D))=t_{1}\text{id}_{V}$ is homothety, it must fix all the $V_{k}$. Hence, the only monomials with nonzero coefficient must satisfy $2n-m\equiv 0\text{ mod }N$. As $n,m\in\{0,\ldots,N\}$, we have $(m,n)$ is $(N,0)$ or $(N,N)$. As $T_{N}(\sqrt{v_{1}v_{2}}(U+D))$ has degree $N$, $T_{N}(\sqrt{v_{1}v_{2}}(U+D))=\sqrt{v_{1}v_{2}}^{N}U^{N}+\sqrt{v_{1}v_{2}}^{N}D^{N}$.
@@
Equations 17 and 18 have no common solutions $x$, so they cannot hold simultaneously. Hence, we must have $d_{0}+d_{1}=0$. ∎
@@
We may think of $d_{0}+d_{1}=0$ and $T_{N}(2-d_{0}^{2})-2=-t_{1}^{2}-t_{2}^{2}-t_{3}^{2}-t_{1}t_{2}t_{3}$ as analogues of the puncture invariant conditions for the skein algebra of the torus by [41, Lem. 7].
@@
Note that if we substitute $d_{0}=A+A^{-1}$ and $d_{1}=-A-A^{-1}$, $P_{k}$ becomes $A^{2}-A^{-2}$, meaning that it corresponds under the homomorphism $\varphi:\mathcal{S}^{RY}(\Sigma_{0,2,2})\to\mathcal{S}(\Sigma_{1,0,0})$ to a nullhomotopic loop in the closed torus. Similarly, if we substitute $d_{0}=A+A^{-1}$ into (19), we recover the condition on the puncture invariant for the once punctured torus at $p=-A^{2}-A^{-2}$, namely, $T_{N}(-A^{2}-A^{-2})=-t_{1}t_{2}t_{3}-t_{1}^{2}-t_{2}^{2}-t_{3}^{3}+2$, given in [41, Lem. 7]. Our investigation was motivated by whether all of the representations of $\mathcal{S}^{RY}(\Sigma_{0,2,2})$ were constructed this way, which we found to the negative.
Problem 6.14.
@@ 7.1 A geometric basis for $S^{RY}(\Sigma_{0,2,2})$
First, consider the case where $\left[\begin{smallmatrix}n_{1}\\ k_{1}\end{smallmatrix}\right]_{g}$ and $\left[\begin{smallmatrix}n_{2}\\ k_{2}\end{smallmatrix}\right]_{g}$ are both loops. We represent them as loops in the four-punctured sphere, respectively $(\frac{n_{1}+k_{1}}{2},n_{1})$ and $(\frac{n_{2}+k_{2}}{2},n_{2})$. Observe we must have $\text{gcd}(\frac{n_{i}+k_{i}}{2},n_{i})=1$, and recall by [3, Lemma 4.1] (see also Section 2.2.5 of [21]) that the geometric intersection number of loops in the four-punctured sphere $(N_{1},K_{1})$ and $(N_{2},K_{2})$ is $2|N_{1}K_{2}-N_{2}K_{1}|$ when $\text{gcd}(N_{i},K_{i})=1$. Applying this when $N_{i}=\frac{n_{i}+k_{i}}{2}$ and $K_{i}=n_{i}$ gives the desired equality.
Now suppose one of the curves, say $\left[\begin{smallmatrix}n_{1}\\ k_{1}\end{smallmatrix}\right]_{g}:=\theta$ is an arc, and the other $\left[\begin{smallmatrix}n_{2}\\ k_{2}\end{smallmatrix}\right]_{g}:=\omega$ is a loop, then let $\theta^{\prime}=v_{1}v_{2}\left[\begin{smallmatrix}n_{1}\\ k_{1}\end{smallmatrix}\right]_{g}^{2}-2=\left[\begin{smallmatrix}2n_{1}\\ 2k_{1}\end{smallmatrix}\right]_{g}$, which is the boundary of a regular neighborhood of $\theta$, as in Example 3.2. Then, from Figure 4 (Left), we see the intersection number of $\omega$ and $\theta^{\prime}$ is twice that of $\theta$ and $\omega$, and, as $\omega$ and $\theta^{\prime}$ are loops. One can show that if $\left[\begin{smallmatrix}n_{1}\\ k_{1}\end{smallmatrix}\right]_{g}$ is an arc then $\text{gcd}(n_{1}+k_{1},2n_{1})=1$, so we can apply [3, Lemma 4.1] as before.
@@ 7.2 Description of pre-image of torus knots
We consider only the case where $0<K<N$, as the other case is similar. By [22, Lemma 1], for any $(N,K)$ such that $N\geq 3$, $0<K<N$ and $\text{gcd}(N,K)=1$, there exist $u,v,w,z$ such that $uz-vw=\pm 1$, $u+w=N$, $v+z=K$, $w\in\{1,\ldots,N-1\}$, $u\in\{1,\ldots,N-2\}$, and $v,z\in\mathbb{Z}_{>0}$. It follows that the pairs $(u,v),(w,z),$ and (uw,zv)(u-w,z-v) are each relatively prime.
@@ 7.3 Structural constants for infinite families of products
To simplify the formulas, we use normalized Chebyshev polynomials $\overline{T}_{n}(x)$, defined by $\overline{T}_{0}(x)=1$ and for all $n\geq 1$, $\overline{T}_{n}(x)=T_{n}(x)$.
@@
$b(p,k)$ $=\begin{cases}A^{-k}\left(\delta_{[p-k]}^{-1}\frac{p-k+1}{4}+\displaystyle{\sum_{h=1}^{\frac{p-k+[p-k]-2}{4}}}([p-k-4h+2]_{A}+[p-k-4h]_{A})\right)&k\leq p-3\quad\text{and}\quad k\equiv p-3\mod 2\\ 0&\text{otherwise}\end{cases}$